X gm of a certain oxide of nitrogen was slowly passed over heated copper to yield CuO and N2 gas. The increase in the weight of copper was 5.33 gm  4 litre of N2 gas at 140C and 754 mm of Hg was also obtained. What is the formula of the oxide.

Grade:

1 Answers

Sunil Kumar FP
9 years ago
we have ,
let the formula of the compound be NaOb,we have to find the formula of the compound.
NaOb +2bCu --------a/2 N2 +bCu2O
mole of N2 formed from ideal gas equation=PV/RT
P=754/760=.99 atm
V=4lt
n= to calculate
T=413 K
n=.99*4/(.0821*413)
n=.117 mole
Thus,a/2=.117,a=.234
difference in weight=143b-63.5*2b=5.33
b=.33
b/a=3/2
Thus the formula is N2O3